OpenStudy (anonymous):

HELP please /: 8. Simplify the following expression. x − 6y + 2z − 5x − 3y + 8z A) –4x + 9y + 10z B) –4x − 9y + 10z C)–4x + 9y − 10z D)4x –9y + 10z 9. Simplify the following expression. 2x − 8y + 3x^2 + 7y − 12x A)3x^2 − 14x − y B)3x^2 − 10x − y C)3x^2 − 10x − 15y D)3x^2 + 10x − y 10. Evaluate the following expression using the values given. Find 4x − 2y if x = −1 and y = −3. A)–10 B)2 C)10 D)–2

jimthompson5910 (jim_thompson5910):

x - 5x = ???

jimthompson5910 (jim_thompson5910):

what do you get when you combine x and -5x

OpenStudy (anonymous):

x-5x?

jimthompson5910 (jim_thompson5910):

1 - 5 = ???

OpenStudy (anonymous):

-4

jimthompson5910 (jim_thompson5910):

so x - 5x or 1x - 5x = -4x

jimthompson5910 (jim_thompson5910):

do the same with the other like terms

OpenStudy (anonymous):

That doesn't make since to me... is there a easier way to do it?

OpenStudy (anonymous):

So WAIT. #8. could be B?

jimthompson5910 (jim_thompson5910):

you basically add/subtract the numbers in front of the variable

jimthompson5910 (jim_thompson5910):

# 8 is choice B

jimthompson5910 (jim_thompson5910):

congrats

jimthompson5910 (jim_thompson5910):

you'll be doing the same thing for #9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

When I am done can you check it for me?

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

#9 is C

jimthompson5910 (jim_thompson5910):

# 9 is not C

jimthompson5910 (jim_thompson5910):

notice how -8y + 7y = -y

OpenStudy (anonymous):

DONT TELL ME THE ANSWER!

jimthompson5910 (jim_thompson5910):

i won't

OpenStudy (anonymous):

there is no -8?

OpenStudy (anonymous):

-8y?

jimthompson5910 (jim_thompson5910):

there's a − 8y

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

yes there is. but not a -8y? it's a subtraction not a negative.

jimthompson5910 (jim_thompson5910):

minus 8y = negative 8y

jimthompson5910 (jim_thompson5910):

they're the same thing

OpenStudy (anonymous):

You acted like it was a negative number.. lol. I don't know :) I am dumb. anyways.

jimthompson5910 (jim_thompson5910):

nah, no you're not

jimthompson5910 (jim_thompson5910):

sry I should have been more clear

OpenStudy (anonymous):

so A is out the question because it can't go down to 14 and then with a variable of y? because I still have 10x.

OpenStudy (anonymous):

So my choices are B & D? Right.

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

I am going to say B.

jimthompson5910 (jim_thompson5910):

B is correct

OpenStudy (anonymous):

YAY :) I get it!!!!

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

#10 is also B

jimthompson5910 (jim_thompson5910):

also correct

jimthompson5910 (jim_thompson5910):

by sheer coincidence, they're all B...

OpenStudy (anonymous):

Yay <3 lol ahahaha.

jimthompson5910 (jim_thompson5910):

congrats lol

OpenStudy (anonymous):

12. x · x = ? ^ this would equal 2x right?

jimthompson5910 (jim_thompson5910):

x times x = x^2, by definition

jimthompson5910 (jim_thompson5910):

x+x = 2x by definition

OpenStudy (anonymous):

so that was right?

jimthompson5910 (jim_thompson5910):

sry i got distracted for a second

jimthompson5910 (jim_thompson5910):

what was right? x*x = 2x ? no, x+x = 2x x times x = x^2

OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −3(2u − 1) –5u − 3 –6u + 3 –5u + 3 –6u − 3 Pemdas? first way I tried it. 2u-1= 1 * -3 =-3? second way I tried it? -3*2-1=-7?

OpenStudy (anonymous):

and 12 is x^2?

jimthompson5910 (jim_thompson5910):

yes, #12 is x^2

jimthompson5910 (jim_thompson5910):

-3 times 2u = ??

OpenStudy (anonymous):

-6?

jimthompson5910 (jim_thompson5910):

-6u

jimthompson5910 (jim_thompson5910):

since there's a 'u' with the 2

jimthompson5910 (jim_thompson5910):

so -3 times 2u = -6u

OpenStudy (anonymous):

so it's -6u and what? and did you know this whole test is nothing but answer B's? I just realized that :D I thought earlier when you said it was all b's it was the past 3 questions but no it's the whole thing. haaha

jimthompson5910 (jim_thompson5910):

very very odd

jimthompson5910 (jim_thompson5910):

-3 times -1 = ??

OpenStudy (anonymous):

and I am on ? number 15.. hahaha! it's weird.

OpenStudy (anonymous):

and 3

OpenStudy (anonymous):

but what's the sign ?

jimthompson5910 (jim_thompson5910):

it's positive since the two negatives cancel

jimthompson5910 (jim_thompson5910):

−3(2u − 1) -3(2u) -3(-1) -6u + 3

OpenStudy (anonymous):

hahahah. that one's b too xD

jimthompson5910 (jim_thompson5910):

that's B..yootiful...lol but very very odd (and concerning)

jimthompson5910 (jim_thompson5910):

but I have verified everything so far, so it's all correct

OpenStudy (anonymous):

this can't be reality! lol

jimthompson5910 (jim_thompson5910):

yeah it's probably a very weird glitch or lazy professors who wrote this assignment

OpenStudy (anonymous):

lmfao. I am sorry. I find this amusing.

jimthompson5910 (jim_thompson5910):

lol me too, wish it didn't do this though

OpenStudy (anonymous):

why? : )

OpenStudy (anonymous):

it;s not like I ain't working out every problem.

jimthompson5910 (jim_thompson5910):

well it just seems odd and it makes me second guess myself

OpenStudy (anonymous):

my dad checked over our work and he is a engineer. he giggled and said that they were all right :) you don't understand how much I appreciate your attendance.

jimthompson5910 (jim_thompson5910):

thanks for the extra verification, I appreciate it

jimthompson5910 (jim_thompson5910):

let's just hope there's a few C's or D's in there lol